# A new paradigm of mortality modeling via individual vitality dynamics

### v2

@@ C.3 Merton’s Problem with Vitality
Appendix D Approximation Methods
D.1 Tangent Approximation
If we set $J(t)=0$, and given the initial vitality level $V(0)=v$, then Jennen and Lerche, (1981); Skiadas and Skiadas, (2014) discussed a tangent approximation method to get the density of $\tau(v)$; that is,
@@ D.2 Series Expansion
Appendix E Cause-of-Death Analysis
@@
Substituting the above equation into equation (E) yields
@@
Summing up equation (E) over $n$ leads to completes the proof of equation (13).
@@
for $n\geq 1$, where the third equation follows from interchanging the order of integrations, while the last equation is derived by changing of variable (letting $\frac{\ln\left(bc^{x}+z\ln c\right)-\ln b}{\ln c}-x=t$). Summing up equation (E) over $n$ ($n\geq 1$) and noting that
$\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau_{Y}}\mathbf{1}\left(\tau_{Y}<\tau_{J,}N(\tau_{Y})=0\right)\right]=\mathrm{e}^{-(q+\lambda)t_{v}^{*}}$

### v3

@@ C.3 Merton’s Problem with Vitality
Definition C.1.
(Admissible Strategies) The pair of strategies $(\pi(\cdot),\zeta(\cdot))$ is called an admissible strategy, if it satisfies the following conditions:
  1. 1.
    $\pi(\cdot)$ and $\zeta(\cdot)$ are progressively measurable;
  2. 2.
    $\mathbb{E}_{0,a,v}\left[\int_{0}^{t}\pi^{2}(s)+\zeta^{2}(s)\mathop{}\!\mathrm{d}s\right]<\infty$ and $A(t)>0$ for any $t\in[0,\infty)$ and $v>0$.
  3. 3.
    associated with $(\pi(\cdot),\zeta(\cdot))$, the state equation $A(\cdot)$ has a unique strong solution.
We denote the set of all admissible strategies by $\Pi$.
@@
Appendix D Disability with Recovery
Proof.
The $T$-year recovery probability for a disabled person aged $x$ can be expressed as
[…]
$=$ $\int_{0}^{\psi}\frac{\Pr\left(\left\{v-Y(T)-W(T)>\psi\right\}\text{ and }\left\{v-Y(t)-W(t)>0\text{ for any }t\in[0,T]\right\}\right)}{F_{0}(\psi)}\mathop{}\!\mathrm{d}F_{0}(v).$ (15)
By assumption, the trend component is $Y(t)=\delta\cdot t$, and we may define
$\widehat{W}(t)$ $=\frac{1}{\sigma}\cdot\left(W(t)+Y(t)\right)=\frac{1}{\sigma}\cdot\left(W(t)+\delta\cdot t\right),$
$\widehat{M}(T)$ $=\frac{1}{\sigma}\cdot\max_{0\leq t\leq T}\left\{W(t)+Y(t)\right\}=\max_{0\leq t\leq T}\left\{\widehat{W}(t)\right\},$
where $\widehat{W}(t)$ is a Brownian motion with drift $\delta/\sigma$, and $\sigma\cdot\widehat{M}(t)$ tracks the maximum vitality depletion until $t$. If $\sigma\cdot\widehat{M}(t)$ is larger than the initial vitality level, then the individual is no longer alive at time $t$. Denote $f(m,w;T)$ as the joint density of $(\widehat{M}(T),\widehat{W}(T))$. Using results from Shreve et al., (2004), one can derive
$f(m,w;T)=\frac{2(2m-w)}{T\sqrt{2\pi T}}e^{\frac{\delta w}{\sigma}-\frac{\delta^{2}T}{2\sigma^{2}}-\frac{(2m-w)^{2}}{2T}},w\leq m,m\geq 0,$
and is zero otherwise. Notice that the numerator of the expression inside equation (15) can be written as a function of $(\widehat{M}(T),\widehat{W}(T))$:
$\Pr\left(\left\{v-Y(T)-W(T)>\psi\right\}\text{ and }\left\{v-Y(t)-W(t)>0\text{ for any }t\in[0,T]\right\}\right)$
$=$ $\Pr\left(v-\sigma\cdot\widehat{W}(T)>\psi,v-\sigma\cdot\widehat{M}(t)>0\right)$
$=$ $\Pr\left(\widehat{W}(T)<\frac{v-\psi}{\sigma},\ \widehat{M}(T)<\frac{v}{\sigma}\right).$
Therefore, substitute the density function $f(m,v;T)$, we have the recovery probability as:
$\int_{0}^{\psi}\frac{1}{F_{0}(\psi)}\cdot\left(\int_{0}^{\frac{v}{\sigma}}\int_{0}^{\frac{v-\psi}{\sigma}}f(m,w;T)\mathop{}\!\mathrm{d}w\mathop{}\!\mathrm{d}m\right)\mathop{}\!\mathrm{d}F_{0}(v).$
A similarly procedure follows for the disability probability and therefore we omit the details.
Appendix E Approximation Methods
E.1 Tangent Approximation
If we set $J(t)=0$, and given the initial vitality level $V(0)=v$, then Jennen and Lerche, (1981) discussed a tangent approximation method to get the density of $\tau(v)$; that is,
@@ E.2 Series Expansion
Appendix F Cause-of-Death Analysis
@@
Substituting the above equation into equation (F) yields
@@
Summing up equation (F) over $n$ leads to completes the proof of equation (13).
@@
for $n\geq 1$, where the third equation follows from interchanging the order of integrations, while the last equation is derived by changing of variable (letting $\frac{\ln\left(bc^{x}+z\ln c\right)-\ln b}{\ln c}-x=t$). Summing up equation (F) over $n$ ($n\geq 1$) and noting that
$\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau_{Y}}\mathbb{1}\left(\tau_{Y}<\tau_{J,}N(\tau_{Y})=0\right)\right]=\mathrm{e}^{-(q+\lambda)t_{v}^{*}}$